OpenStudy (anonymous):

Tell whether the ordered pair is a solution of the given system. (3,1) x+3y=6 4x-5y=7 _____________ x+3y=6 4x-5y=7

OpenStudy (anonymous):

substitute x = 3 and y = 1 in both equations and check LHS = x+3y = (3) + 3(1) = 3+3 = 6 = RHS LHS = 4x - 5y = 4(3) - 5(1) = 12 - 5 = 7 = RHS Since x=3 and y=1 satisfy both the equations, hence it is a solution to the system